# Enzo Erlich Université Paris Cité, CNRS, IRIF, F-75013, Paris, France and EPITA Research Laboratory (LRE), Paris, France

### v3

@@ 1 Introduction
More recently, another notion of limited nondeterminism has received considerable attention: history-deterministic automata [11, 25] [^1] There is a closely related notion, good-for-gameness, which is often, but not always equivalent [3] (despite frequently being used interchangeably in the past). are nondeterministic automata whose nondeterminism can be resolved based on the run constructed thus far, but independently of the remainder of the input. This property makes history-deterministic automata suitable for composition with games, trees, and other automata, applications which classically require deterministic automata. History-determinism has been studied in the context of regular [1, 25, 30], pushdown [23, 31], quantitative [4, 11], and timed automata [24]. For automata that can be determinized, history-determinism offers the potential for succinctness (e.g., co-Büchi automata [30]) while for automata that cannot be determinized, it even offers the potential for increased expressiveness (e.g., pushdown automata [23, 31]). In the quantitative setting, the exact power of history-determinism depends largely on the type of quantitative automata under consideration. So far, it has been studied for quantitative automata in which runs accumulate weights into a value using a value function such as Sum, LimInf, Average, and that assign to a word the supremum among the values of its runs. For these automata, history-determinism turns out to have interesting applications for quantitative synthesis [3]. Here, we continue this line of work by investigating history-deterministic Parikh automata, a mildly quantitative form of automata.
@@
Despite their expressiveness, nondeterministic Parikh automata retain some decidability: nonemptiness, in particular, is NP-complete [15] and regular separability is decidable [10]. For weakly unambiguous Parikh automata inclusion [8] is decidable. On the other hand, for deterministic Parikh automata, finiteness, universality, regularity, and model-checking are decidable as well [6, 29]. Figueira and Libkin [15] also argued that Parikh automata are well-suited for querying graph databases, while mitigating some of the complexity issues related with more expressive query languages. Further, they have been used in the model checking of transducer properties [17].
@@
Secondly, we show that history-deterministic Parikh automata satisfy almost the same closure properties as deterministic ones, the only difference being non-closure under complementation. This result has to be contrasted with unambiguous Parikh automata being closed under complement [7]. Thus, history-determinism is a too strong form of nondeterminism to preserve closure under complementation, a phenomenon that has already been observed in the case of pushdown automata [23, 31].
Finally, we study the algorithmic properties of history-deterministic Parikh automata. Most importantly, safety model checking for HDPA is decidable, as it is for PA. The problem asks, given a system and a set of bad prefixes specified by an automaton, whether the system has an execution that has a bad prefix. This allows, for example, to check properties of an arbiter of some shared resource like “the accumulated waiting time between requests and responses of client 1 is always at most twice the accumulated waiting time for client 2 and vice versa”. Note that this property is not $\omega$-regular.
@@ 2 Definitions
An alphabet is a finite nonempty set $\Sigma$ of letters. As usual, $\varepsilon$ denotes the empty word, $\Sigma^{*}$ denotes the set of finite words over $\Sigma$, $\Sigma^{+}$ denotes the set of finite nonempty words over $\Sigma$, and $\Sigma^{\omega}$ denotes the set of infinite words over $\Sigma$. The length of a finite word $w$ is denoted by $|w|$ and, for notational convenience, we define $|w|=\infty$ for all infinite words $w$. Finally, $|w|_{a}$ denotes the number of occurrences of the letter $a$ in a finite word $w$.
@@ 2.1 Semilinear Sets
Proposition 1 ([19]).
If $C,C^{\prime}\subseteq\mathbb{N}^{d}$ are semilinear, then so are $C\cup C^{\prime}$, $C\cap C^{\prime}$, $\mathbb{N}^{d}\setminus C$, as well as $\mathbb{N}^{d^{\prime}}\cdot C$ and $C\cdot\mathbb{N}^{d^{\prime}}$ for every $d^{\prime}\geqslant 1$.
@@ 2.3 Parikh Automata
The Parikh automaton $(\mathcal{A},C)$ is deterministic if for every state $q$ of $\mathcal{A}$ and every $a\in\Sigma$, there is at most one pair $(q^{\prime},\vec{v})\in Q\times D$ such that $(q,(a,\vec{v}),q^{\prime})$ is a transition of $\mathcal{A}$. Note that this definition does not coincide with $\mathcal{A}$ being a DFA: As mentioned above, $\mathcal{A}$ accepts words over $\Sigma\times D$ while $(\mathcal{A},C)$ accepts words over $\Sigma$. Therefore, determinism is defined with respect to $\Sigma$ only.
Note that the above definition of $L(\mathcal{A},C)$ coincides with the following alternative definition via accepting runs: A run $\rho$ of $(\mathcal{A},C)$ is a run
@@ 3 History-deterministic Parikh Automata
Figure 2: The automaton for Example 3
@@ 4 Expressiveness
We begin by stating and proving a pumping lemma for HDPA. The basic property used here, just as for the pumping lemmata for PA and DPA [6], is that shuffling around cycles of a run does not change whether it is accepting or not, as acceptance only depends on the last state of the run being accepting and the vectors (and their multiplicity) that appear on the run, but not the order of their appearance.
@@
starting and ending in the same state and such that the $q_{j}$ are pairwise different. Now, let $p$ be the number of states of $\mathcal{A}$ and let $m$ be the number of cycles of $\mathcal{A}$. Note that every run infix containing at least $p$ transitions contains a cycle.
@@
It is instructive to compare our pumping lemma for HDPA to those for PA and DPA [6]:
@@
First, we consider the separation between DPA and HDPA. The language $N$ from Example 3, which is accepted by an HDPA, is known to be not accepted by any DPA: DPA are closed under complementation [29] while the complement of $N$ is not even accepted by any PA [7].
To show that PA are more expressive than HDPA, consider the language $E=\{a,b\}^{*}\cdot\{a^{n}b^{n}\mid n>0\}$, which can easily be seen to be accepted by a PA. We show that $E$ is not accepted by any HDPA  [^2] Note that the related language $\{a,b\}^{*}\cdot\{a^{n}\#a^{n}\mid n\in\mathbb{N}\}$ is not accepted by any DPA [6]. via an application of the pumping lemma.
@@
Cadilhac et al. studied unambiguity in Parikh automata in the guise of unambiguous constrained automata (UCA) [7]. Constrained automata are a related model and effectively equivalent to PA. Intuitively, an UCA $(\mathcal{A},C)$ over an alphabet $\Sigma$ consists of an unambiguous $\varepsilon$-NFA $\mathcal{A}$ over $\Sigma$, say with $d$ transitions, and a semilinear set $C\subseteq\mathbb{N}^{d}$, i.e., $C$ has one dimension for each transition in $\mathcal{A}$. It accepts a word $w\in\Sigma^{*}$ if $\mathcal{A}$ has an accepting run processing $w$ (due to unambiguity this run must be unique) such that the Parikh image of the run (recording the number of times each transition occurs in the run) is in $C$.
On the other hand, Bostan et al. introduced so-called weakly-unambiguous Parikh automata (WUPA) [5]. Intuitively, a WUPA $(\mathcal{A},C)$ over $\Sigma$ is a classical PA as introduced here where every input over $\Sigma$ has at most one accepting run (in the sense of the definition in Section 2). Bostan et al. discuss the different definitions of unambiguity and in particular show that every UCA is a WUPA, but that WUPA are strictly more expressive [5]. Here, we compare the expressiveness of HDPA to that of UCA and WUPA.
@@
The language $E$ from the proof of Theorem 1 is accepted by an UCA [7] and thus also by a WUPA, as every UCA can be turned into an equivalent WUPA [5]. However, $E$ is not accepted by any HDPA, as shown in that proof. This yields the first two separations.
@@
So, consider some $w\in E_{b}$, i.e., $w=w^{\prime}ba^{n}b^{n}$ for some $n>0$ and some $w^{\prime}\in\{a,b\}^{*}$. Define $m=|w^{\prime}b|$ and note that $c^{m}w$ is in $E^{\prime}$, i.e., accepted by $(\mathcal{A},C)$. Hence, let $\rho$ be the unique run of $(\mathcal{A},C)$ processing $c^{m}w$, let $\rho^{\prime}$ be the prefix of $\rho$ processing $c^{m}$, and let $\rho^{\prime\prime}$ be the suffix processing $w$. So, there is some $q$ (the state $\rho^{\prime}$ ends in, which is equal to the state $\rho^{\prime\prime}$ begins with) such that $\rho^{\prime}\in R_{q}$ and the extended Parikh image $\vec{v}\,^{\prime}$ induced by $\rho^{\prime}$ is in $C_{q}$. Also, $\rho^{\prime\prime}$ is an run of $\mathcal{A}_{q}$ starting in $q$ and ending in $F$. Let $\vec{v}$ be the extended Parikh image induced by $\rho^{\prime\prime}$. Note that $\vec{v}+\vec{v}\,^{\prime}$ is the extended Parikh image induced by the full run $\rho=\rho^{\prime}\rho^{\prime\prime}$ that witnesses $c^{m}w\in L(\mathcal{A},C)$, and is therefore in $C$. From this we conclude $\vec{v}\in C_{q}^{\prime}$ and therefore that $\rho^{\prime\prime}$ is an accepting run of $(\mathcal{A}_{q},C_{q}^{\prime})$ processing $w$, i.e., $w\in L(\mathcal{A}_{q},C_{q}^{\prime})$ and $R_{q}$ is nonempty as witnessed by $\rho^{\prime}$.
For the other direction, consider a $w\in L(\mathcal{A}_{q},C_{q}^{\prime})$ for some $q$ with nonempty $R_{q}$. Then there is an accepting run $\rho^{\prime\prime}$ of $(\mathcal{A},C_{q}^{\prime})$ processing $w$, say with induced extended Parikh image $\vec{v}\in C_{q}^{\prime}$. By construction, there is also a $\vec{v}\,^{\prime}\in C_{q}$ such that $\vec{v}+\vec{v}\,^{\prime}\in C$. Furthermore, $\vec{v^{\prime}}$ is the extended Parikh image induced by some run $\rho^{\prime}$ of $(\mathcal{A},C)$ processing some word of the form $c^{m}$. Now, $\rho^{\prime}\rho^{\prime\prime}$ is a run of $(\mathcal{A},C)$ starting in the initial state, processing $c^{m}w$, ending in $F$, and with extended Parikh image $\vec{v}+\vec{v}\,^{\prime}\in C$, i.e., it is an accepting run. Hence, $c^{m}w\in L(\mathcal{A},C)=E^{\prime}$. As $w$ does not contain any $c$ ($(\mathcal{A}_{q},C_{q}^{\prime})$ has no $c$-transitions), this implies that $w$ must be of the form $w^{\prime}ba^{n}b^{n}$ with $w^{\prime}\in\{a,b\}^{m-1}$, i.e., $w\in E_{b}$ as required.
@@
which is a variation of the language $N$ of words that have a non-Dyck prefix. In $N^{\prime}$, the length of that prefix is given by the number of $c$’s in the beginning of the word, which makes accepting the language easier. Nevertheless, Bostan et al. showed that $N^{\prime}$ is not accepted by any UCA, but by a WUPA [5].
@@
3.) The (disjoint) union $T\cup E$ is accepted by some PA, as both $T$ and $E$ are accepted by a PA and PA are closed under union. But it is neither accepted by any HDPA nor by any WUPA, as both models are closed under intersection [^5] For WUPA, this was shown by Bostan et al. [5], for HDPA this is shown in Theorem 5 on Page 5., i.e., $(T\cup E)\cap\{a,b\}^{*}=E$ and $(T\cup E)\cap\{c,d\}^{*}=T$ yield the desired separations.
@@ 4.2 History-deterministic Reversal-bounded Counter Machines
There is one more automaton model that is closely related to Parikh automata, i.e., reversal-bounded counter machines, originally introduced by Ibarra [27]. These are, in their most general form, two-way automata with multiple counters that can be incremented, decremented, and tested for zero, but there is a fixed bound on the number of reversals of the reading head and on the number of switches between increments and decrements (on each counter). It is known that Parikh automata and nondeterministic reversal-bounded counter machines are equivalent [29], while deterministic reversal-bounded counter machines are strictly more expressive than deterministic Parikh automata [6]. Here, we compare history-deterministic reversal-bounded counter machines and history-deterministic Parikh automata (and, for technical reasons, also history-deterministic Parikh automata with $\varepsilon$-transitions).
@@
The problem with counter machines is that even their deterministic variants are Turing-complete [27]. Therefore, one must impose some restrictions in order to obtain decidability of some decision problems. One way to do that is to introduce bounds on the number of reversals of the direction the reading head moves and on the number of reversals of the stack height of each counter.
@@
    Every RBCM can be effectively turned into an equivalent 1-RBCM.
@@
    Every RBCM can be effectively turned into an equivalent one where the number of reversals of each counter is bounded by $1$. This construction preserves determinism and one-wayness.
@@
    1-DRBCM are strictly more expressive than DPA.
@@
2.) We first show how to turn an HDPA into an equivalent 1-RBCM and then show that it is history-deterministic. This construction is inspired by the constructions turning a PA (a DPA) into an equivalent 1-RBCM (an equivalent 1-DRBCM) [6, 29].
Intuitively, the 1-RBCM simulates the given HDPA by keeping track of the extended Parikh image of a run in its counters, i.e., there is a bijection between the runs of the HDPA and the 1-RBCM. As the HDPA is one-way by definition, so is the simulating RBCM. When the end of the input word is reached, acceptance depends on whether the extended Parikh image (the counter values) are in $C$. The 1-RBCM can do so using the following result due to Cadilhac et al. [6]: membership in a given semilinear set can be tested by a DRBCM that does not move its reading head. More formally they showed that for every semilinear set $C\subseteq\mathbb{N}^{d}$, there exists a DRBCM $\mathcal{M}_{C}$ such that, for any configuration $S_{0}=(q_{I},\rhd w\lhd,i,\vec{v}\cdot\vec{0})$ with $\vec{v}\in\mathbb{N}^{d}$, the unique run starting from $S_{0}$ is accepting if and only if $\vec{v}\in C$. Moreover, $\mathcal{M}_{C}$ does not move its reading head. Note that $\mathcal{M}_{C}$ will in general use more than $d$ counters. In the configuration $S_{0}$, the first $d$ counters contain the vector to be checked, and the remaining ones are initialized with zero. This is exactly the situation after the simulation of the HDPA has finished.
It remains to argue that the resulting 1-DRBCM is indeed history-deterministic. However, this is straightforward, since it simulates the history-deterministic PA (recall that there is a bijection between the runs, which can be used to transfer the resolver) and then deterministically checks membership in a semilinear set.
@@
Step 3) Finally, we need to show how to eliminate $\varepsilon$-transitions in $\varepsilon$-PA while preserving history-determinism. The construction is similar to one used by Klaedtke and Rueß [28] to show that Parikh automata are closed under homomorphisms, which also requires the removal of $\varepsilon$-transitions. For the sake of completeness, we present the construction here, as we need to show that it allows us to preserve history-determinism.
@@
is such that $b_{j}$ is equal to $1$ if $K_{j}$ has been removed at least once from $\rho$ to obtain $\mathrm{red}(\rho)$. Otherwise, $b_{j}$ is equal to 00. Note that this results in finitely many transitions, as there are only finitely many reduced $\varepsilon$-$a$-$\varepsilon$-paths and only finitely many choices for the $\chi(\rho)$. [^6] Furthermore, it is not hard to see that the construction can be made effective, as one does not have to consider all $\varepsilon$-$a$-$\varepsilon$-path: every transition in $\mathcal{A}^{\prime}$ is witnessed by an $\varepsilon$-$a$-$\varepsilon$-path of bounded length.

### v4

@@ 1 Introduction
More recently, another notion of limited nondeterminism has received considerable attention: history-deterministic automata [13, 28] [^1] There is a closely related notion, good-for-gameness, which is often, but not always equivalent [4] (despite frequently being used interchangeably in the past). are nondeterministic automata whose nondeterminism can be resolved based on the run constructed thus far, but independently of the remainder of the input. This property makes history-deterministic automata suitable for composition with games, trees, and other automata, applications which classically require deterministic automata. History-determinism has been studied in the context of regular [1, 28, 33], pushdown [26, 34], quantitative [5, 13], and timed automata [27] (see [6] for a recent survey). For automata that can be determinized, history-determinism offers the potential for succinctness (e.g., co-Büchi automata [33]) while for automata that cannot be determinized, it even offers the potential for increased expressiveness (e.g., pushdown automata [26, 34]). In the quantitative setting, the exact power of history-determinism depends largely on the type of quantitative automata under consideration. So far, it has been studied for quantitative automata in which runs accumulate weights into a value using a value function such as Sum, LimInf, Average, and that assign to a word the supremum among the values of its runs. For these automata, history-determinism turns out to have interesting applications for quantitative synthesis [4]. Here, we continue this line of work by investigating history-deterministic Parikh automata. These are a mildly quantitative form of automata with a significant gap between the expressiveness and algorithmic properties of its deterministic and nondeterministic versions, making it a good candidate for studying restricted forms of nondeterminism, such as history-determinism.
@@
Despite their expressiveness, nondeterministic Parikh automata retain some decidability: nonemptiness, in particular, is NP-complete [18] and regular separability is decidable [12]. For weakly unambiguous Parikh automata, inclusion is decidable [10]. On the other hand, for deterministic Parikh automata, finiteness, universality, regularity, and model-checking are decidable as well [8, 32]. Figueira and Libkin [18] also argued that Parikh automata are well-suited for querying graph databases, while mitigating some of the complexity issues related with more expressive query languages. Further, they have been used in the model checking of transducer properties [20].
@@
Secondly, we show that history-deterministic Parikh automata satisfy almost the same closure properties as deterministic ones, the only difference being non-closure under complementation. This result has to be contrasted with unambiguous Parikh automata being closed under complement [9, Proposition 5]. Thus, history-determinism is a too strong form of nondeterminism to preserve closure under complementation, a phenomenon that has already been observed in the case of pushdown automata [26, 34].
Further, we study the algorithmic properties of history-deterministic Parikh automata. Most importantly, safety model checking for HDPA is decidable, as it is for PA. The problem asks, given a system and a set of bad prefixes specified by an automaton, whether the system has an execution that has a bad prefix. This allows, for example, to check properties of an arbiter of some shared resource like “the accumulated waiting time between requests and responses of client 1 is always at most twice the accumulated waiting time for client 2 and vice versa”. Note that this property is not $\omega$-regular.
@@
This article is a revised and extended version of an article published at CONCUR 2023 [17], which contains all proofs omitted in the conference version and a new section on resolvers implementable by Parikh automata.
@@ 2 Definitions
An alphabet is a finite nonempty set $\Sigma$ of letters. As usual, $\varepsilon$ denotes the empty word, $\Sigma^{*}$ denotes the set of finite words over $\Sigma$, and $\Sigma^{+}$ denotes the set of finite nonempty words over $\Sigma$. The length of a finite word $w$ is denoted by $|w|$ and, for notational convenience, we define $|w|=\infty$ for all infinite words $w$. Finally, $|w|_{a}$ denotes the number of occurrences of the letter $a$ in a finite word $w$.
@@ 2.1 Semilinear Sets
Ginsburg and Spanier showed that semilinear sets are closed under the Boolean operators [22] while closure under concatenation can easily be shown by extending the generating vectors with zeroes.
Proposition 1.
  1. 1.
    If $C,C^{\prime}\subseteq\mathbb{N}^{d}$ are semilinear, then so are $C\cup C^{\prime}$, $C\cap C^{\prime}$, $\mathbb{N}^{d}\setminus C$.
  2. 2.
    If $C\subseteq\mathbb{N}^{d}$ and $C^{\prime}\subseteq\mathbb{N}^{d^{\prime}}$ are semilinear, then so is $C\cdot C^{\prime}$.
@@ 2.3 Parikh Automata
The Parikh automaton $(\mathcal{A},C)$ is deterministic if for every state $q$ of $\mathcal{A}$ and every $a\in\Sigma$, there is at most one pair $(q^{\prime},\vec{v})\in Q\times D$ such that $(q,(a,\vec{v}),q^{\prime})$ is a transition of $\mathcal{A}$. Note that this definition does not coincide with $\mathcal{A}$ being a DFA: As mentioned above, $\mathcal{A}$ accepts words over $\Sigma\times D$ while $(\mathcal{A},C)$ accepts words over $\Sigma$. Therefore, determinism is defined with respect to $\Sigma$ only. Further, $(\mathcal{A},C)$ is complete if for every $q\in Q$ and every $a\in\Sigma$, there is a $\vec{v}\in D$ and a $q^{\prime}\in Q$ such that $(q,(a,\vec{v}),q^{\prime})\in\Delta$, i.e., from every state, every letter can be processed. As usual, an incomplete automaton can be completed by adding a fresh non-accepting sink state.
Note that the above definition of $L(\mathcal{A},C)$ coincides with the following alternative definition via accepting runs.
Definition 1.
A run $\rho$ of $(\mathcal{A},C)$ is a run
@@
Note that there is a closely related automaton model, called constrained automata [8], that has the same expressive power as PA as introduce here [8, Theorem 3.4.]. Here, we mostly work with PA, but we compare unambiguous constrained automata in Subsection 4.1 to history-deterministic PA.
@@ 3 History-deterministic Parikh Automata
Figure 2: The automaton for Example 3.
@@ 4 Expressiveness
We begin by stating and proving a pumping lemma for HDPA. The basic property used here, just as for the pumping lemmata for PA and DPA [8, Lemma 3.5 and Lemma 3.6], is that shuffling around cycles of a run does not change whether it is accepting or not, as acceptance only depends on the last state of the run being accepting and the vectors (and their multiplicity) that appear on the run, but not the order of their appearance.
@@
starting and ending in the same state and such that the $q_{j}$ for $j\in\{0,1,\ldots,n\}$ are all pairwise different. Now, let $p$ be the number of states of $\mathcal{A}$ and let $m$ be the number of cycles of $\mathcal{A}$. Note that every run infix containing at least $p$ transitions contains a cycle.
@@
It is instructive to compare our pumping lemma for HDPA to those for PA and DPA [8, Lemma 3.5 and Lemma 3.6]:
@@
First, we consider the separation between DPA and HDPA. The language $N$ from Example 3, which is accepted by an HDPA, is known to be not accepted by any DPA: DPA are closed under complementation [32, Property 4] while the complement of $N$ is not even accepted by any PA [9, Proposition 11].
To show that PA are more expressive than HDPA, consider the language
$E=\{a,b\}^{*}\cdot\{a^{n}b^{n}\mid n>0\},$
which can easily be seen to be accepted by a PA. We show that $E$ is not accepted by any HDPA  [^2] Note that the related language $\{a,b\}^{*}\cdot\{a^{n}\#a^{n}\mid n\in\mathbb{N}\}$ is not accepted by any DPA [8, Proposition 3.8]. via an application of the pumping lemma.
@@
Cadilhac et al. studied unambiguity in Parikh automata in the guise of unambiguous constrained automata (UCA) [9] (recall that (possibly ambiguous) constrained automata are effectively equivalent to PA [8, Theorem 3.4]). Intuitively, an UCA $(\mathcal{A},C)$ over an alphabet $\Sigma$ consists of an unambiguous $\varepsilon$-NFA $\mathcal{A}$ over $\Sigma$, say with $d$ transitions, and a semilinear set $C\subseteq\mathbb{N}^{d}$, i.e., $C$ has one dimension for each transition in $\mathcal{A}$. It accepts a word $w\in\Sigma^{*}$ if $\mathcal{A}$ has an accepting run processing $w$ (due to unambiguity this run must be unique) such that the Parikh image of the run (recording the number of times each transition occurs in the run) is in $C$.
On the other hand, Bostan et al. introduced so-called weakly-unambiguous Parikh automata (WUPA) [7]. Intuitively, a WUPA $(\mathcal{A},C)$ over $\Sigma$ is a classical PA as introduced here where every input over $\Sigma$ has at most one accepting run processing it (in the sense of Definition 1), i.e., the vectors labelling the transitions are ignored as described in that definition. Bostan et al. discuss the different definitions of unambiguity and in particular show that every UCA is a WUPA, but that WUPA are strictly more expressive (see Remark 11 of [7] and the references therein). Here, we compare the expressiveness of HDPA to that of UCA and WUPA.
@@
The language $E$ from the proof of Theorem 1 is accepted by an UCA [9, Theorem 12] and thus also by a WUPA, as every UCA can be turned into an equivalent WUPA (again, see Remark 11 of [7]). However, $E$ is not accepted by any HDPA, as shown in the proof of Theorem 1. This yields the first two separations.
@@
So, consider some $w\in E_{b}$, i.e., $w=w^{\prime}ba^{n}b^{n}$ for some $n>0$ and some $w^{\prime}\in\{a,b\}^{*}$. Define $m=|w^{\prime}b|$ and note that $c^{m}w$ is in $E^{\prime}$, i.e., accepted by $(\mathcal{A},C)$. Hence, let $\rho$ be the unique run of $(\mathcal{A},C)$ processing $c^{m}w$, let $\rho^{\prime}$ be the prefix of $\rho$ processing $c^{m}$, and let $\rho^{\prime\prime}$ be the suffix processing $w$. So, there is some $q$ (the state $\rho^{\prime}$ ends in, which is equal to the state $\rho^{\prime\prime}$ begins with) such that $\rho^{\prime}\in R_{q}$ and the extended Parikh image $\vec{v}\,^{\prime}$ induced by $\rho^{\prime}$ is in $C_{q}$. Also, $\rho^{\prime\prime}$ is a run of $\mathcal{A}_{q}$ starting in $q$ and ending in $F$. Let $\vec{v}$ be the extended Parikh image induced by $\rho^{\prime\prime}$. Note that $\vec{v}+\vec{v}\,^{\prime}$ is the extended Parikh image induced by the full run $\rho=\rho^{\prime}\rho^{\prime\prime}$ that witnesses $c^{m}w\in L(\mathcal{A},C)$, and is therefore in $C$. From this we conclude $\vec{v}\in C_{q}^{\prime}$ and therefore that $\rho^{\prime\prime}$ is an accepting run of $(\mathcal{A}_{q},C_{q}^{\prime})$ processing $w$, i.e., $w\in L(\mathcal{A}_{q},C_{q}^{\prime})$ and $R_{q}$ is nonempty as witnessed by $\rho^{\prime}$.
For the other direction, consider a $w\in L(\mathcal{A}_{q},C_{q}^{\prime})$ for some $q$ with nonempty $R_{q}$. Then there is an accepting run $\rho^{\prime\prime}$ of $(\mathcal{A}_{q},C_{q}^{\prime})$ processing $w$, say with induced extended Parikh image $\vec{v}\in C_{q}^{\prime}$. By construction, there is also a $\vec{v}\,^{\prime}\in C_{q}$ such that $\vec{v}+\vec{v}\,^{\prime}\in C$. Furthermore, $\vec{v^{\prime}}$ is the extended Parikh image induced by some run $\rho^{\prime}$ of $(\mathcal{A},C)$ processing some word of the form $c^{m}$. Now, $\rho^{\prime}\rho^{\prime\prime}$ is a run of $(\mathcal{A},C)$ starting in the initial state, processing $c^{m}w$, ending in $F$, and with extended Parikh image $\vec{v}+\vec{v}\,^{\prime}\in C$, i.e., it is an accepting run. Hence, $c^{m}w\in L(\mathcal{A},C)=E^{\prime}$. As $w$ does not contain any $c$ ($(\mathcal{A}_{q},C_{q}^{\prime})$ has no $c$-transitions), this implies that $w$ must be of the form $w^{\prime}ba^{n}b^{n}$ with $w^{\prime}\in\{a,b\}^{m-1}$, i.e., $w\in E_{b}$ as required.
@@
which is a variation of the language $N$ of words that have a non-Dyck prefix. In $N^{\prime}$, the length of that prefix is given by the number of $c$’s in the beginning of the word, which makes accepting the language easier. Nevertheless, Cadilhac et al. showed that $N^{\prime}$ is not accepted by any UCA [9, Proposition 14], but Bostan et al. showed that it is accepted by a WUPA [7, Remark 11].
@@
3.) The (disjoint) union $T\cup E$ is accepted by some PA, as both $T$ and $E$ are accepted by a PA and PA are closed under union. But it is neither accepted by any HDPA nor by any WUPA, as both models are closed under intersection [^5] For WUPA, this was claimed by Bostan et al. [7], for HDPA this is shown in Theorem 5 on Page 5., i.e., $(T\cup E)\cap\{a,b\}^{*}=E$ and $(T\cup E)\cap\{c,d\}^{*}=T$ yield the desired separations.
@@ 4.2 History-deterministic Reversal-bounded Counter Machines
There is one more automaton model that is closely related to Parikh automata, i.e., reversal-bounded counter machines, originally introduced by Ibarra [30]. These are, in their most general form, two-way automata with multiple counters that can be incremented, decremented, and tested for zero, but there is a fixed bound on the number of reversals of the reading head and on the number of switches between increments and decrements (on each counter). It is known that Parikh automata and nondeterministic reversal-bounded counter machines are equivalent [31, Property 8], while deterministic reversal-bounded counter machines are strictly more expressive than deterministic Parikh automata [8, Proposition 3.14]. Here, we compare history-deterministic reversal-bounded counter machines and history-deterministic Parikh automata (and, for technical reasons, also history-deterministic Parikh automata with $\varepsilon$-transitions).
@@
The problem with counter machines is that even their deterministic variants are Turing-complete [30, Theorem 3.3]. Therefore, one must impose some restrictions in order to obtain decidability of some decision problems. One way to do that is to introduce bounds on the number of reversals of the direction the reading head moves and on the number of reversals of the stack height of each counter.
@@
    Every RBCM can be effectively turned into an equivalent 1-RBCM [30, Theorem 2.2].
@@
    Every RBCM can be effectively turned into an equivalent one where the number of reversals of each counter is bounded by $1$. This construction preserves determinism and one-wayness [2, Theorem 5].
@@
    1-DRBCM are strictly more expressive than DPA [8, Proposition 3.14].
@@
2.) We first show how to turn an HDPA into an equivalent 1-RBCM and then show that it is history-deterministic. This construction is inspired by the constructions turning a PA into an equivalent 1-RBCM [32, Property 8] and turning a DPA into an equivalent 1-DRBCM [8, Proposition 3.14].
Given an HDPA $(\mathcal{A},C)$, the simulating counter machine $\mathcal{M}$ works in two phases. In the first phase, $\mathcal{M}$ simulates the run of $(\mathcal{A},C)$ (using the same states as $(\mathcal{A},C)$) and computes the extended Parikh image of the simulated run using its counters. Hence, in the first phase, there is a bijection between the runs of $(\mathcal{A},C)$ and $\mathcal{M}$. If the simulated run does not end in an accepting state, then the run of $\mathcal{M}$ ends in a rejecting state after the first phase. However, if the run ends in an accepting state, then the second phase begins.
In this case, acceptance of $(\mathcal{A},C)$ depends on whether the extended Parikh image of the run (stored in the counter values of $\mathcal{M}$) is in $C$. We can do so using the following result due to Cadilhac et al. [8, Proposition 3.13]: membership in a given semilinear set can be tested by a DRBCM that does not move its reading head. More formally they showed that for every semilinear set $C^{\prime}\subseteq\mathbb{N}^{d}$, there exists a DRBCM $\mathcal{M}_{C^{\prime}}$ such that, for any configuration $S_{0}=(q_{I},\rhd w\lhd,i,\vec{v}\cdot\vec{0})$ with $\vec{v}\in\mathbb{N}^{d}$, the unique run of $\mathcal{M}_{C^{\prime}}$ starting from $S_{0}$ is accepting if and only if $\vec{v}\in C^{\prime}$. Moreover, $\mathcal{M}_{C^{\prime}}$ does not move its reading head. Note that $\mathcal{M}_{C^{\prime}}$ will in general use more than $d$ counters. In the configuration $S_{0}$, the first $d$ counters contain the vector to be checked, and the remaining ones are initialized with zero. This is exactly the situation after the first phase of the simulation of the HDPA described above has finished. Thus, in the second phase, we run DRBCM $\mathcal{M}_{C}$ to check whether the extended Parikh image of the simulated run of $(\mathcal{A},C)$ stored in the counter values of the simulating run of $\mathcal{M}$ is in $C$.
As there is a bijection between runs of $(\mathcal{A},C)$ and the runs of $\mathcal{M}$ restricted to their first phase, it is straightforward to argue that $\mathcal{M}$ accepts $L(\mathcal{A},C)$. Furthermore, $\mathcal{M}$ is one-way, as $(\mathcal{A},C)$ is one-way by definition, which implies that the first phase of the simulation is one-way as well, while the head does not move in the second phase.
It remains to argue that the resulting 1-RBCM $\mathcal{M}$ is indeed history-deterministic. However, this is straightforward, since it simulates the history-deterministic PA $(\mathcal{A},C)$ (recall that there is a bijection between the runs, which can be used to transfer the resolver) and then deterministically checks membership in a semilinear set. Thus, using the bijection, a resolver for $(\mathcal{A},C)$ can directly be turned into a resolver for $\mathcal{M}$.
@@
Step 3) Finally, we need to show how to eliminate $\varepsilon$-transitions in $\varepsilon$-PA while preserving history-determinism. The construction is similar to one used by Klaedtke and Rueß [31, Theorem 20] to show that Parikh automata are closed under homomorphisms, which also requires the removal of $\varepsilon$-transitions. For the sake of completeness, we present the construction here, as we need to show that it allows us to preserve history-determinism.
@@
is such that $b_{j}$ is equal to $1$ if $K_{j}$ has been removed at least once from $\rho$ to obtain $\mathrm{red}(\rho)$. Otherwise, $b_{j}$ is equal to 00. Note that this results in finitely many transitions, as there are only finitely many reduced $\varepsilon$-$a$-$\varepsilon$-paths and only finitely many choices for the $\chi(\rho)$. [^6] Furthermore, it is not hard to see that the construction can be made effective, as one does not have to consider all $\varepsilon$-$a$-$\varepsilon$-paths: every transition in $\mathcal{A}^{\prime}$ is witnessed by an $\varepsilon$-$a$-$\varepsilon$-path of bounded length.